### v1

# Cost-Efficient Long Code Translation using LLMs while Leveraging Identifier Replacements

Abstract
In the domain of software development, LLMs have been utilized to automate tasks such as code translation, where source code from one programming language is translated to another while preserving its functionality. However, LLMs often struggle with long source codes that don’t fit into the context window, which produces inaccurate translations. To address this, we propose a novel zero-shot code translation method that incorporates identifier replacement. By substituting user-given long identifiers with generalized placeholders during translation, our method allows the LLM to focus on the logical structure of the code, by reducing token count and memory usage, which improves the efficiency and cost-effectiveness of long code translation. Our empirical results demonstrate that our approach preserves syntactical and hierarchical information and produces translation results with reduced tokens.
1 Introduction
Large Language Models (LLMs) demonstrate impressive abilities in capturing text semantics abstractly, without the need for task-specific training Radford et al. (2018); Brown et al. (2020); Arora et al. (2023); Weidinger et al. (2022). This characteristic of “general intelligence" enables LLMs to function as zero-shot predictors across a variety of downstream tasks, including question answering Li et al. (2023), document retrieval Pradeep et al. (2023), and code translation Pan et al. (2024a). However, the practical application of LLMs in software engineering faces several significant challenges. One notable area is code translation, which involves converting large, complex codebases from one programming language to another. This process is essential for organizations aiming to migrate legacy systems or to integrate various programming languages within large code repositoriesAhmad et al. (2023); Pan et al. (2024b); Krishna et al. (2021); Nitin et al. (2022). It is also critical to ensure that the functional and logical integrity of the code is maintained throughout the translation processAhmad et al. (2023); Pan et al. (2024b). But the inherent context length limit of every LLM poses a significant challenge for code translation tasks Radford et al. (2018); Brown et al. (2020). Long code bases frequently contain intricate dependencies, and this limitation often necessitates partitioning the code into smaller segments. This division complicates the translation process and increases the risk of losing structural consistency across different sections of the code Jana et al. (2023); Gong et al. (2024).
Source code identifiers, particularly complex user-defined function names, class names, and variable names, play a crucial role in structuring large codebases. These identifiers often contain descriptive metadata or project-specific conventions, leading to increase token usage when processed by LLMs. While identifiers improve code readability for human developers, it can contribute to rapid token consumption in LLMs models, reducing the available context window for performing code translation. To alleviate the above challenges, we present a novel zero-shot code translation method that employs an identifier replacement  [^1] https://anonymous.4open.science/r/LCT-LLM/ strategy to optimize the translation process using LLMs. While token reduction is the most immediate benefit of our identifier replacement strategy, a subtler but equally important reason for improved accuracy is the shift in the model’s attention distribution. Long and descriptive identifiers often dilute the attention of LLMs across semantically redundant sub-tokens. By compressing these identifiers into compact placeholders (e.g., id_1), the model is nudged to prioritize the underlying syntactic and control-flow structure of the program rather than overfitting to identifier surface forms. This mechanism effectively reduces noise and enables the LLM to better capture execution-relevant relationships (e.g., function calls, data dependencies) during translation. In other words, identifier replacement transforms the problem into a more syntax-driven translation task, which explains the observed accuracy improvements for procedural languages in our experiments. This tailored extraction process refines the preprocessing stage by differentiating between critical identifiers, such as function and class names, and non-essential ones, thereby enhancing translation fidelity. By simplifying the task for the LLM and prioritizing the logical structure of the code over lengthy identifiers, our approach enables the model to manage longer code sequences within its token limit more effectively Jana et al. (2023)  Dinh et al. (2024). This not only decreases the number of tokens needed to represent the code but also significantly reduces the computational cost of inference.
Our Contributions:
To summarize, the following are our contributions in this work.
  1. 1.
    We present a novel zero-shot code translation method that replaces lengthy, complex user-defined identifiers with shorter, generalized placeholders. This approach significantly reduces token count and memory usage, resulting in a more efficient and cost-effective solution for translating long code using LLMs.
  2. 2.
    We provide empirical evidence showing that our proposed solution effectively manages long code sequences while preserving both syntactical and hierarchical information during the translation task.
2 Related Work
Since our work mainly investigates deep metric learning based long code translation approaches, we now first discuss the recent development in machine translation literature, and then follow it up with how these models are adapted particularly for code translation, a task which we, in fact, address in this paper.
Generic Machine Translation:
Machine translation (MT) has been an active research area for decades, evolving from rule-based approaches to modern deep learning-based systems. Traditional statistical MT models Och (2003) used phrase-based techniques, but they struggled with maintaining long-range dependencies. The advent of Neural Machine Translation (NMT) Bahdanau (2014) introduced attention mechanisms, significantly improving translation quality. Later, the Transformer model Vaswani et al. (2017) revolutionized MT by efficiently handling long sequences through self-attention. Recent works have extended NMT for various domains, including document-level translation (Maruf et al., 2021), where maintaining global coherence is critical. However, these models often suffer from “context truncation issues” when handling lengthy documents, a limitation similar to what we observe in long code translation. Addressing this, recent research has explored context-aware translation (Voita et al., 2018), yet these approaches remain underexplored in code-related tasks.
Code Translation:
Automated code translation is essential for software portability and legacy system modernization. Early approaches relied on rule-based translation, which required extensive manual effort. With the rise of machine learning, statistical methods (Koehn et al., 2003) and deep learning-based techniques (Post and Junczys-Dowmunt, 2024) have been employed to improve translation accuracy. Recent works have leveraged the pretrained knowledge of language for code translation, considering it as a sequence-to-sequence task. Models such as TransCoder Roziere et al. (2020), CodeT5 Wang et al. (2021), and CodeBERT (Feng et al., 2020) have demonstrated strong performance in multilingual code translation. Furthermore, execution-based evaluation has been proposed as an alternative to BLEU scores, ensuring functional correctness (Kulal et al., 2019). Despite these advancements, LLMs face severe context-length limitations when dealing with large codebases. This has prompted research into chunking strategies (Dikert et al., 2016) and hierarchical decoding (Zhou et al., 2022) to improve translation efficiency. However, these techniques often introduce fragmentation, causing loss of cross-file dependencies in large code repositories.
Long Code Translation:
Long code translation presents unique challenges, particularly for LLMs constrained by finite context windows. Recent research has explored hierarchical attention mechanisms (Zhou et al., 2022), external memory augmentation (Wang et al., 2024). Although these methods improve efficiency, they often require specialized fine-tuning or additional memory resources, making them computationally expensive. Some works focus on segmenting code intelligently using AST-based chunking (Lin et al., 2021), but this still does not resolve the problem of ‘identifier fragmentation’, where long identifiers are split across multiple chunks, degrading model performance. Alternative methods like ‘syntax-guided translation’ (Liu et al., ) attempt to retain structural consistency, yet those frameworks fail to optimize memory usage effectively. A promising direction in this space is token reduction techniques. For instance, Pan et al. (2024c) proposed a prompt compression strategy for reducing token count in LLM-based translation. However, such methods primarily focus on natural language tasks and are not optimized for structured code translation.
Unlike prior chunking-based LCT methods that introduce fragmentation, we propose a zero-shot identifier replacement strategy that directly reduces token count without altering code semantics. While existing models require fine-tuning or additional memory resources, our method is model-agnostic and can be applied to any pretrained LLM. By dynamically substituting long identifiers with compact placeholders, our approach significantly reduces computational cost, making it ideal for industry-scale code translation tasks. To the best of our knowledge, this is the first study to introduce an identifier-aware preprocessing framework for cost-efficient long code translation.
3 Methodology
Input: Code file name, programming language, optional dataset name, optional target directory
Output: Mapping of original identifiers to new identifiers, modified code
1
2
3
4Function Identifier_Extraction(code, language):
5   LANGUAGE $\leftarrow$ GetLanguage(languagelanguage) $\triangleright$ Retrieve programming language metadata
6   parser $\leftarrow$ GetParser(languagelanguage);
7 $\triangleright$ Initialize a syntax parser
8   tree $\leftarrow$ parser.parse(Encode(codecode)) $\triangleright$ Source code -> Syntax tree
9   root_node $\leftarrow$ tree.root_node;
10 $\triangleright$ Extract root node from syntax tree
11
12  grammar $\leftarrow$ LoadGrammar(language);
13 $\triangleright$ Load predefined grammar
14   acceptable_parents $\leftarrow$ grammar[‘PARENT’];
15 $\triangleright$ Retrieve valid parent nodes for identifiers
16   unacceptable_parents $\leftarrow$ grammar[‘NOT’];
17 $\triangleright$ Retrieve invalid parent nodes for filtering
18
19  query $\leftarrow$ LANGUAGE.query(‘‘(identifier) @identifier");
20 $\triangleright$ Query to extract identifiers
21   qc $\leftarrow$ query.captures(root_node);
22 $\triangleright$ Get matching identifiers
23   captures $\leftarrow$ [ ];
24 $\triangleright$ Initialize an empty list for extracted identifiers
25
26 foreach capture in qc do
27      node $\leftarrow$ capture[0];
28    $\triangleright$ Extract identifier node
29      parent $\leftarrow$ node.parent;
30    $\triangleright$ Extract the parent node of the identifier
31
32    if parent.type not in acceptable_parents or code[node.start_byte:node.end_byte] in unacceptable_parents then
33       continue;
34       $\triangleright$ Skip if parent invalid or identifier predefined
35    Append(captures, capture);
36    $\triangleright$ Store valid identifiers
37 return captures;
38 $\triangleright$ Return the list of extracted identifiers
39
40Function GetIdSet(code, captures):
41   id_set $\leftarrow$ CreateSet();
42 $\triangleright$ Initialize an empty set for unique identifiers foreach capture in captures do
43      node $\leftarrow$ capture[0];
44    $\triangleright$ Extract identifier node
45    Add(id_set, code[node.start_byte:node.end_byte]) $\triangleright$ Add to set
46 return id_set;
47 $\triangleright$ Return the set of unique identifiers
48
49Function Process_Code(file_name, language, dataset, target_dir):
50   code $\leftarrow$ LoadData(file_name, dataset);
51 $\triangleright$ Load source code from file
52   language $\leftarrow$ GetLang(language);
53 $\triangleright$ Retrieve language metadata
54  id_list $\leftarrow$ Identifier_Extraction(code, language);
55 $\triangleright$ Extract identifiers from the code
56   identifiers $\leftarrow$ GetIdSet(code, id_list);
57 $\triangleright$ Retrieve unique identifiers
58  i $\leftarrow 0$;
59 $\triangleright$ Initialize index counter
60   mapping $\leftarrow$ {};
61 $\triangleright$ Initialize mapping dictionary
62   len_reduced $\leftarrow$ 0;
63 $\triangleright$ Initialize length reduction counter
64 foreach identifier in identifiers do
65      candidate $\leftarrow$ id_i;
66    $\triangleright$ Generate shorter placeholder for identifier
67    if TOKEN_Length(candidate) $\geq$ TOKEN_Length(identifier) then
68         mapping[identifier] $\leftarrow$ identifier;
69       $\triangleright$ Keep original identifier if replacement is not shorter
70    else
71         len_reduced $\leftarrow$ len_reduced + TOKEN_Length(identifier) - TOKEN_Length(candidate);
72         mapping[identifier] $\leftarrow$ candidate;
73       $\triangleright$ Replace identifier with a shorter placeholder
74         i $\leftarrow$ i + 1;
75       $\triangleright$ Increment placeholder index
76
77
78 foreach (old, new) in mapping do
79    if old in [‘‘file_name", ‘‘language", ‘‘len_reduced", ‘‘dataset"] then
80       continue;
81       $\triangleright$ Skip protected keywords from replacement
82     code $\leftarrow$ Replace(code, old, new);
83    $\triangleright$ Replace original identifier with its mapped value
84 return mapping, code;
85 $\triangleright$ Return updated mapping and modified code
Algorithm 1 Process Code and Replace Identifiers
Figure 1: An overview of the identifier extraction and replacement algorithm for long source code translation. The process involves identifier extraction, classification into syntactic categories, replacement using an identifier mapping strategy, code translation, and final restoration to ensure syntactic and semantic correctness.
In this section, we outline the task description and provide a detailed explanation of our proposed approach as depicted in the Figure 1, to perform long code translation with LLMs through the use of identifier replacement.
Task Description:
The task of translating source code in a zero-shot scenario with a LLM $\mathcal{M}$ can be described as using a multinomial sampling strategy to convert code written in a source language $\mathcal{L}_{s}$ into corresponding code in a target language $\mathcal{L}_{t}$. Formally speaking, given an input code sequence $C_{s}=\{x_{1},x_{2},\ldots,x_{n}\}$ in $\mathcal{L}_{s}$, the $\mathcal{M}$ generates the output sequence $C_{t}=\{y_{1},y_{2},\ldots,y_{m}\}$ in $\mathcal{L}_{t}$, such that the functionality of $C_{s}$ is preserved.
To mitigate context length limitations, we utilize an identifier extraction and replacement strategy during translation. First, we identify the set of unique long identifiers, $\mathcal{I}_{s}=\{i_{1},i_{2},\dots,i_{k}\}$, within the source code $C_{s}$. These identifiers are then replaced with a set of generalized placeholders, $\mathcal{P}=\{p_{1},p_{2},\ldots,p_{k}\}$, resulting in a simplified and shorter source sequence $C^{\prime}_{s}=\{x^{\prime}_{1},x^{\prime}_{2},\dots,x^{\prime}_{n}\}$. The LLM translates $C^{\prime}_{s}$ in the target language $\mathcal{L}_{t}$. Finally, in a post-processing step, the placeholders identifiers $\mathcal{P}$ are mapped back to corresponding original identifiers $\mathcal{I}_{s}$ within the translated code $C_{t}$, preserving the code’s functionality and semantics.
Identifier Extraction and Replacement:
We utilize tree-sitter Brunsfeld and contributors , a robust parsing tool, to systematically extract relevant long identifiers from source code. By generating accurate syntax trees across various programming languages, tree-sitter enables us to identify key programming constructs, which are then considered as primary candidates [^2] A detailed statistics of the probable identifier list has been reported in Table 4 for identifier replacements. Following are the key programming constructs serve as primary candidates for identifier extraction and replacement.
  • Function Definitions: To capture function names and associated parameter identifiers.
  • Variable Declarations: To extract variable names, ensuring the preservation of variable semantics across the translation process.
  • Class and Method Declarations: To capture class names and method signatures, crucial for maintaining object-oriented structure.
  • Reserved Identifiers: Identifiers such as this, self, super, null, true, and false are not replaced to maintain language-specific semantics. Additionally, built-in function names such as print (Python), System.out.println (Java), std::cout (C++), and keywords like def, class, static, and return are preserved to ensure syntactic correctness.
Employing these key programming constructs, we developed grammars for each programming language which are then used to extract the potential identifiers suitable for replacements. Let $\mathcal{I}_{s}=\{i_{1},i_{2},\ldots,i_{k}\}$ represent the set of extracted identifiers. For each $i_{j}\in\mathcal{I}_{s}$, we assign a unique placeholder $p_{j}\in\mathcal{P}$ of a small token length = $2$ and produce a simplified source code $C^{\prime}_{s}$ from original source code $C_{s}$. This way, we achieve the length reduction, $\Delta l$, in the token length as given by: $\Delta l=\sum_{j=1}^{k}\left(|i_{j}|-|p_{j}|\right)$, where $|i_{j}|$ and $|p_{j}|$ denote the token length of the original identifier and its corresponding placeholder, respectively. After translation is complete and the LLM generates the output sequence $C^{\prime}_{t}$ in the target language, a reverse process is applied to restore the original identifiers. Let $g$ represent the reverse mapping function, which takes translated sequence $C^{\prime}_{t}$ and original set of identifiers $\mathcal{I}_{t}$ in the target language, and produces the final translated sequence $C_{t}$:
$C_{t}=g(C^{\prime}_{t},\mathcal{I}_{t})=\{y_{1},y_{2},\dots,y_{m}\}$
$\quad g:C^{\prime}_{t}\times\mathcal{I}_{t}\rightarrow C_{t}$
This post-processing step ensures that placeholders in the translated code are replaced with the corresponding identifiers from the target language, preserving the semantics of the original source code. This complete method is described in the Algorithm 1.
4 Experimental Setup
Figure 2: The figure illustrates the transformation of source code during the identifier replacement (IdRep) process. (a) represents the original source code with long, descriptive identifiers, while (b) shows the modified code where identifiers are replaced with shorter placeholders. This transformation helps reduce token count, enabling LLMs to process longer sequences efficiently while preserving semantic correctness.
In this section, we provide an overview of the dataset used in our experiments. We then explain various closed and open source LLMs, along with their specific configurations, to assess their capacity in translating long code sequences using identifier replacements.
4.1 Dataset
Original Dataset Reduced Dataset Length C C++ Java Go Python C C++ Java Go Python 2000 714 90,015 568 51,418 6,596 20 1,717 577 33 37 4000 214 19,545 70 11,812 2,221 7 416 109 4 17 8000 22 4,134 3 2,048 258 1 162 22 0 4
Table 1: Statistics of XcodeEval dataset. The table shows the total number of code instances across five programming languages (C, C++, Java, Go, Python) in both the Original and Reduced subsets.
To analyze our approach of code translation via identifier replacement, we partitioned the dataset into three subsets based on token length: those with lengths greater than or equal to 2000, 4000, and 8000 tokens. These lengths were determined using the TikToken tokenizer Pan et al. (2024c). We also filtered out instances that lacked corresponding test cases to ensure accurate evaluation. Table 1 represents the total number of source code instances across various languages for both the original and reduced subsets. The “Original Dataset" provide raw counts prior to filtering, while the “Reduced Dataset" columns illustrate the instances that were retained.
4.2 LLM Models
In this work, we use models such as GPT3.5-Turbo Ye et al. (2023), GPT4o-Mini Hurst et al. (2024), Mixtral (8*7B) Jiang et al. (2023) [^3] It is noteworthy that, while CodeLlama and Mixtral models are theoretically capable of handling inputs more than the mentioned tokens, they often produce noise beyond this limit., and CodeLlama Grattafiori et al. (2023). The context window of these models during inference are 16K, 128K, 8K and 4K respectively. The model choice is driven by their widespread use and popularity, as well as the constraints of computing capabilities.
5 Results & Analysis
Language # of Samples GPT3.5-Turbo GPT 4o-Mini Mixtral 8*7B CodeLlama NoRep IdRep NoRep IdRep NoRep IdRep NoRep IdRep 2000 C 20 37.1% 37.5% 41.0% 45.7% 10.3% 11.0% 12.1% 13.0% C++ 1717 10.5% 10.4% 17.0% 14.5% 2.6% 2.6% 3.6% 3.3% Java 577 14.3% 15.0% 21.3% 20.5% 4.3% 3.6% 3.9% 4.1% Go 33 35.5% 38.0% 58.5% 61.3% 7.6% 8.0% 8.3% 8.8% Python 37 32.8% 35.0% 33.3% 40.0% 8.6% 9.0% 9.3% 10.3% 4000 C 7 22.5% 24.5% 41.0% 41.3% 5.8% 5.4% - - C++ 416 14.0% 16.0% 25.0% 20.3% 2.8% 1.4% - - Java 109 11.8% 10.9% 17.5% 14.5% 2.3% 2.0% - - Go 4 - - - - - - - - Python 17 20.3% 28.6% 32.8% 35.0% 5.3% 5.8% - - 8000 C 1 - - - - - - - - C++ 162 7.5% 9.8% 15.8% 16.5% - - - - Java 22 7.4% 6.8% 16.0% 17.8% - - - - Go 0 - - - - - - - - Python 4 - - - - - - - -
Table 2: Performance comparison on the XcodeEval dataset across the LLMs, under two configurations: No Replacement (NoRep) and Identifier Replacement (IdRep). The table is organized by sample size (2000, 4000, and 8000) and five programming languages (C, C++, Java, Go, and Python). Performance is evaluated using accuracy metric. The final performance is computed over the average for each source language. The best-performing results for each setting are highlighted in bold. Due to insufficient samples (less than 5 instances), we do not report the average accuracy for certain languages. Additionally, due to context length limitations, we do not report results for Mixtral and CodeLlama for the 4000 and 8000 token length buckets.
In this section, we present the experimental results and significant observations organized around key research questions (RQ) related to long code translations using identifier replacements. To evaluate our identifier replacement approach (IdRep) for long code translation, we establish a baseline using translations without identifier replacement (NoRep). Additionally, an example comparing the IdRep and NoRep methods is depicted in Figure 2. A translation is deemed successful only if it compiles, executes, and passes all unit tests provided by a particular dataset. Our evaluation indeed covers both syntactic correctness and functional behavior. We therefore emphasize that our evaluation validates functionality under the available test suite, but broader runtime behavior remains an open challenge for future work. Thus, in this work we investigate the following key research questions.
  • RQ-1: What is the impact of identifier replacements on translation performance for long code across different programming languages, LLM models, and different token length buckets?
  • RQ-2: To what extent do identifier replacements reduce token length in LLM-based code translation across different programming languages?
Language Token Saved (per sample) C 22.76 Java 11.34 Go 17.21 C++ 5.75 Python 15.29
Table 3: This table presents average number of tokens saved per sample across different programming languages in the XcodeEval dataset when applying the IdRep strategy. The reduction in token count enables more efficient utilization of LLM context windows, benefiting languages with extensive identifier usage.
Identifier replacement vs. language:
The translation accuracies with and without identifier replacements (IdRep and NoRep, respectively) for various programming languages, token length buckets and LLMs are presented in the Table 2. We observed that the Identifier replacement, as a context length reduction strategy during translation, tends to be more effective for procedural programming languages (e.g., C, Go) than object-oriented languages (e.g., C++, Java). For instance, Identifier replacement in the 2000-token bucket yields contrasting results for GPT-4o model where translation accuracy improves significantly for C (41.0% to 45.7%), while the performance decreases for C++ (17.0% to 14.5%). This can be due to relatively simpler structure of procedural programming languages compared to object-oriented languages, where meaning of an identifier is highly context-dependent. By substituting these identifiers with abstract placeholders, the model risks losing vital contextual information necessary for maintaining code correctness, which could potentially disrupt object-oriented dependencies.
Table 3 illustrates the efficacy of identifier replacement in reducing token counts across various programming languages, enabling larger code segments to be accommodated within the LLM context window for translation. Thus, replacing identifiers can significantly reduce translation costs for long code sequences when using LLMs. This is particularly beneficial for translating complex industry-specific codebases that often use lengthy, user-defined identifiers.
Identifier replacement vs. LLMs:
In the Table  2, we observed an inconsistency in improvements across models which indicates that identifier replacement benefits the code translation performance of larger models (e.g., GPT) more than smaller models (e.g., Mixtral, CodeLlama). This suggests weaker abstraction in smaller models and a greater reliance on explicit semantic cues for accurate long code translation. Although the average reduction may appear modest in relative terms, it is important to note that this reduction often targets the most critical bottleneck positions long user-defined identifiers. These identifiers not only inflate token counts but also frequently exceed subword segmentation boundaries, leading to poor generalization by LLMs. By systematically replacing such high-cost tokens, our method creates space for including additional lines of logic that would otherwise be truncated. Thus, the benefit is not only in percentage reduction but in strategically freeing context window capacity where it matters most.
6 Ablation Study
Category Code Elements Function (F) compact_constructor_declaration constructor_declaration method_declaration function_declaration function_definition function_declarator Declaration (D) variable_declarator declaration array_declarator var_spec initializer_pair pointer_declarator Class (C) class_definition class_declaration interface_declaration enum_declaration annotation_type_declaration Expression (E) assignment_expression initializer_list assignment parenthesized_expression array_initializer Parameter (P) optional_parameter_declaration typed_parameter parameter_declaration receiver_parameter formal_parameter lambda_parameters default_parameter
Table 4: The categorization of Identifiers in the source Code. This table classifies identifiers into five major categories based on their syntactic roles in source code: Function (F) for function and method definitions, Declaration (D) for variable and construct declarations, Class (C) for class and interface declarations, Expression (E) for assignment and initialization expressions, and Parameter (P) for function and method parameters.
We conducted an ablation study to isolate the impact of identifier semantics on long code translation. We selectively replaced identifiers in different syntactic categories to analyze how this affects translation performance based on the identifiers’ structural role. We categorized identifiers into five major categories (as shown in Table 4): function definitions, variable declarations, class-related elements, method names, and function parameters. These categories represent key elements in program structure and logic. This study used the GPT-4o-Mini model with a 2000-token length bucket to translate Java and C++ codes to other languages.
Effective long code translation through identifier replacement necessitates syntactic awareness specific to each programming language. For instance, when translating from Java, identifier replacement (IdRep) accuracy is 20.5% when all identifiers are replaced. However, restricting replacement to only function or expression identifiers significantly improves accuracy to approximately 22%. Similarly, translating from C++, selectively replacing only parameter identifiers yields the best performance (23.5%). This significantly outperforms both benchmark translations with no identifier replacement (17%) and translations with all the identifiers replaced (14.5%).
Further, our findings show that replacing declaration identifiers harms translation performance across programming languages, likely due to their fundamental role in code structure. In contrast, replacing parameter and expression identifiers significantly improves long code translation when using identifier replacements.
Java C++ Acc. NoRep Acc. IdRep Acc. Acc. NoRep Acc. IdRep Acc. Declaration 19.7% 21.3% 20.5% 16.7% 17.0% 14.5% Function 22.2% 16.7% Parameter 16.0% 23.5% Expression 22.7% 17.7%
Table 5: Performance comparison of Java and C++ translations across different identifier replacement categories. The table evaluates the translation performance when selectively replacing different identifier types (Declaration, Function, Parameter, and Expression) for Java and C++ while translating them into C, C++, Go, and Python. The accuracy (Acc.) represents the mean accuracy across the target languages for each identifier category, while the IdRep Accuracy (IdRep Acc.) represents the average accuracy when all identifier types are replaced and the NoRep Accuracy (NoRep Acc.) represents the average accuracy without any replacement as reported in Table 2. The best performance is in bold faced.
7 Conclusion and Future Directions
We present a novel zero-shot approach for long code translation that leverages identifier replacement to overcome the context window and memory limitations of LLMs. By substituting long identifiers with shorter placeholders, the method reduces token count, leading to cost-effective and efficient translation without sacrificing code integrity. Our experiments demonstrate significant cost savings and improved translation accuracy.
Future research can focus on hybrid techniques, such as combining identifier replacement with symbolic execution, to further boost translation performance.
8 Limitations
While our identifier replacement strategy significantly reduces token length and enhances computational efficiency in long code translation, several limitations remain. First, the approach assumes that replacing long identifiers with placeholders does not affect semantic interpretation, but this assumption may not always hold, particularly for programming languages with strong type systems and deep interdependencies between identifiers. In object-oriented languages like Java and C++, identifier replacement may lead to loss of context, affecting method resolution and dependency tracking. Additionally, the effectiveness of identifier replacement varies across programming paradigms, with procedural languages (e.g., C, Go) benefiting more than those with complex inheritance and polymorphism. Another limitation arises in the availability datasets for long codebases to empirically validate our proposed approach. Finally, our evaluation primarily focuses on syntactic correctness, leaving functional correctness and runtime behavior as open challenges that require further investigation.